OpenStudy (anonymous):

Find a point on the y-axis that is equidistant from (5,-5) an (1,1)

OpenStudy (anonymous):

let the point (0, y) is equidistant from the given points (5 - 0)^2 + (-5 - y)^2 = (1 - 0)^2 + (1 - y)^2 => 25 + 25 + y^2 + 10y = 1 + 1 + y^2 - 2y => 50 + 10y = 2 - 2y => 12y = -48 y = -4 The required point is (0, -4)

OpenStudy (anonymous):

@Athenalexis do you get it?

OpenStudy (anonymous):

@Hannah_Waller yes I have one more question could you help me ?

OpenStudy (anonymous):

i can try @Athenalexis

OpenStudy (anonymous):

@Hannah_Waller Show that the points A(-2, 6), B(5, 3), C(-1, -11) and D(-8, -8) are the vertices of a rectangle.

OpenStudy (anonymous):

IM not real sure how to do this one i looked it up andthis seems to mabye be something like what you are looking for, http://www.homeworklib.com/question_13834_a-rectangle-is-defined-by-the-vertices-a65-b121-c813-and-d107#.VTeyX5PCefg hope it helps

OpenStudy (anonymous):

@Athenalexis

OpenStudy (anonymous):

thank you so much @Hannah_Waller